OpenStudy (anonymous):

The side of an isosceles right triangle of hypotenuse 5sqrt(2) cm is :

OpenStudy (anonymous):

@jiteshmeghwal9 ...cn u help

OpenStudy (jiteshmeghwal9):

Isosceles Triangle has its two sides equal . therefore Let the perpendicular & base be x both. According to Pythagoras Theorem , we have\[\Huge{(p)^2+(b)^2=(h)^2}\]\[\Huge{(x)^2+(x)^2=(5\sqrt{2})^2}\]

OpenStudy (jiteshmeghwal9):

solve this for x @basith :)

OpenStudy (jiteshmeghwal9):

what do u gt @basith ??

OpenStudy (anonymous):

giv me a min

OpenStudy (jiteshmeghwal9):

ok !

OpenStudy (anonymous):

\[2x ^{2}=50\] \[x^{2}=(\frac{ 50 }{ 2 })=25\] am i right???

OpenStudy (jiteshmeghwal9):

ya !

OpenStudy (jiteshmeghwal9):

now, x=?

OpenStudy (jiteshmeghwal9):

Hint:- square root both sides

OpenStudy (anonymous):

5

OpenStudy (jiteshmeghwal9):

yup ! that's ur answer :)

OpenStudy (anonymous):

thnx.....

OpenStudy (jiteshmeghwal9):

yw :)